OpenStudy (anonymous):

simplify e^x(e^-x + 1) - e^-x(e^x + 1)

OpenStudy (anonymous):

= e + e^x - e - e ^-x = e^x - e^-x

OpenStudy (anonymous):

where'd the ones go?